OpenStudy (anonymous):

8x - 3 - 2x = -45 8x-2x=6 6x=-45-3 6x=-41 x=-41/6 x=-6.83 or 6.83

OpenStudy (mimi_x3):

6x = -42 x = -7

OpenStudy (saifoo.khan):

8x - 3 - 2x = -45 8x-2x=6x 6x=-45+3 6x=-42 x=-42/6 x=-7

OpenStudy (anonymous):

need to add 3 to both sides rather than subtract

OpenStudy (anonymous):

what on earth? \[8x-3-2x=-45\] \[6x=-45+3=-42\] \[x=\frac{-42}{6}=-7\]

OpenStudy (anonymous):

oooooooooooh i see this line \[8x-2x=6x\] is an explanation. got it!